OpenStudy (anonymous):

(4y^a - 6x^b)(4y^a + 6x^b)

OpenStudy (anonymous):

same thing

OpenStudy (anonymous):

wow thanks

OpenStudy (campbell_st):

what do you actually need to do... as its not an equation so can't subtract 6x from both sides

OpenStudy (anonymous):

now im confused.. i needed to foil this..

OpenStudy (zzr0ck3r):

yes foild this first. There is no equation so you cant do ^^^^ all that. Do you know how to foil it?

OpenStudy (campbell_st):

its is simply the difference of 2 squares... if you are exanding (a + b)(a - b ) = a^2 - b^2

OpenStudy (anonymous):

uhhm i know how to set it up.. like first outter inner last but i dunno wtf 4ya times 4ya would come out to be..

OpenStudy (campbell_st):

so you will have \[(4y^a)^2 - (6x^b)^2 \] that you need to simplify

OpenStudy (anonymous):

sooo 16y^(a2)-36x^(b2) ??

OpenStudy (zzr0ck3r):

you dont need to actually, use what campell said example: (4x^2 -16b^2) = ((2x)^2 - (4b)^2) = (2x-4)(2x+4)

OpenStudy (anonymous):

thanks erryybody!

OpenStudy (campbell_st):

correct

OpenStudy (campbell_st):

\[16y^{2a} - 36x^{2b}\]